OpenStudy (anonymous):

for the given equation, list the intercepts and test for symmetry. 4x^2+16y^2=64

OpenStudy (anonymous):

if i could get step by step that would be awsome so I will know how to do it please

OpenStudy (anonymous):

??

OpenStudy (rajee_sam):

\[4x ^{2} + 16y ^{2} = 64\] divide the whole thing by 64\[\frac{ 4x ^{2} + 16 y ^{2} }{64} = \frac{ 64 }{ 64 }\] \[\frac{ x ^{2} }{ 16} +\frac{ y ^{2} }{ 4 } = 1\] This is an equation of an Ellipse centered at \[(0,0)\]

OpenStudy (rajee_sam):

\[\frac{ x ^{2} }{ a ^{2} } + \frac{ y ^{2} }{ b ^{2} } = 1\] x is the major axis and y is the minor axis. because \[a > b\]

OpenStudy (rajee_sam):

 You can see the intercepts there